OpenStudy (anonymous):

Write an algebraic equation for the following. The difference of a number squared and the number is one less than four times the number. A. (x-x)^2=4x-1 B. x-x^2=4x-1 C. x^2-x=1-4x D. x^2-x=4x-1

OpenStudy (anonymous):

If your number is x, the first thing you do to it is square it, so now you have x^2 Then it says it is the difference (difference indicates subtraction) between the number squared and itself, so if your number is x, then you have x^2-x The next thing it says is that it is equal to something else x^2-x= Then it says that it is equal to one less than something x^2-x=___-1 Finally, it says that it is one less than four times your number, x x^2-x=4x-1 So your answer is d. x^2-x = 4x-1